OpenStudy (anonymous):

Evaluate r+s/8 r=44 and s = 20 simplify your answer

OpenStudy (anonymous):

r + s/8 44 + 20/8 44 + (5*4)/4*2 44 + 5/2 44 * (2/2) + 5/2 88/2 + 5/2 93/2 (or.. = 46.5)

jhonyy9 (jhonyy9):

44 +20/8 = 44 + 5/2 =(88+5)/2 = 93/2